OpenStudy (anonymous):

Solve the system of equations algebraically. I know i have to change one of these equations to y = mx + b: 2x - y = 6 and -3x + 4y = 1. I just don't know what to do from there.

OpenStudy (anonymous):

do you know how to solve to equation by y ?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

oh wait. i think i know how to do this. do you change the -3x + 4y = 1 to y = mx + b? And then solve for y?

OpenStudy (anonymous):

thats how you solve for y it is when you change it into y = mx + b you want to know what "y" so you are trying to " solve for it" this goes the same as any other variable

OpenStudy (anonymous):

don't you need to solve for x too?

OpenStudy (anonymous):

yes you do it says "solve a system of equations" and you are given x y 2x - y = 6 -3x + 4y = 1. I

OpenStudy (anonymous):

only 2 :D

OpenStudy (anonymous):

first of all put them in slope intercept form

OpenStudy (anonymous):

i did, but i didn't get some clean numbers like integers.

OpenStudy (anonymous):

fractions

OpenStudy (anonymous):

but you know how to solve systems of equations right?

OpenStudy (anonymous):

just sub them back in and see if the are correct

OpenStudy (anonymous):

for fractions just get the reciprocal

OpenStudy (anonymous):

and try to take them out

OpenStudy (anonymous):

Is slope intercept form the only way to go through it? Can you leave them as the orinal ones and get rid of y by multiplying all of the numbers by 4 for 2x - y = 24?

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

solving by graphing

OpenStudy (anonymous):

i believe i am solving the equations algrebraically. But i am not sure what you mean by graphing...

OpenStudy (anonymous):

graphing them and see where they intersect (not the advisable method)

OpenStudy (anonymous):

and there also elimination method

OpenStudy (anonymous):

well another check is by subbing the x and y values and see if it works

OpenStudy (anonymous):

slope intercept is not a way or method it is how the equation must be lined up before any calculations take place

OpenStudy (anonymous):

if they are all put in slope intercept form it should look like this y = 2 x - 6 y = 3x/4 + 1.1/4

OpenStudy (anonymous):

does that look ok? :/

OpenStudy (anonymous):

if i were to use elimination method because you to cancel and "eliminate" one of the terms distributive property is extensively used here multiply -3 [ ( 2 x ) - 6 = y ] multiply 8 [ (3x/4 ) + 1.1/4 = y ]

OpenStudy (anonymous):

until get the solution for y then substitute it to get x